OpenStudy (anonymous):

HELP -2(x+3)=-2(x+1)-4

OpenStudy (precal):

distribute -2

OpenStudy (precal):

-2x-6=-2x-2-4

OpenStudy (precal):

your -2x will cancel out

OpenStudy (precal):

Your solution is all real numbers

OpenStudy (precal):

-2x-6=-2x-6

OpenStudy (precal):

see how both sides are the same

OpenStudy (anonymous):

ya, so is the answer all real numbers

OpenStudy (precal):

yes

OpenStudy (anonymous):

thank you :)

OpenStudy (precal):

because all numbers will work

OpenStudy (precal):

gotta go

OpenStudy (anonymous):

k bye

OpenStudy (mathstudent55):

When the variables cancel out, and you get a true statement, the answer is all numbers. When the variables cancel out, and you get a false statement, the answer is no solution. Here, you have -2x - 6 = -2x - 6 Add 2x to both sides to get -6 = -6 The variables canceled out, and you have a true statement, so the answer is all real numbers.